OpenStudy (anonymous):

Calculate the wavelength of the radiation released when an electron moves from n=5 to n=2

OpenStudy (anonymous):

\[1/\lambda = R ( 1/na ^{2} -1/nb ^{2})\] \[R = 1.097 \times 10^{7}\] \[1/\lambda = R (1/(2)^{2}-1/(5)^{2})\] \[1/\lambda = R (99/25)\] \[\lambda = 25/99R\]

OpenStudy (anonymous):

thank you!